OpenStudy (anonymous):

The line of symmetry of the parabola whose equation is y = ax^2 - 4x + 3 is x = -2. What is the value of "a"? -2 -1 -1/2

OpenStudy (anonymous):

if x=-2 is the axis of symmetry,then its eq. is x+2=0 squaring x^2+4x+4=0 -x^2-4x-4=0 hence a=-1

OpenStudy (anonymous):

So it's -1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks..!

OpenStudy (anonymous):

yw